# Deep Stochastic Kinematic Models for Probabilistic Motion Forecasting in Traffic

### v3

@@ IV-B1 Formulation 1: Velocity Components
Figure 4: Average Displacement Error (ADE) across training across each method and formulation in the small dataset setting. Learning with stochastic kinematic priors aids with faster learning. Compared to the baseline (blue), all models employing kinematic priors converge much more quickly. Formulation 4 (brown) converges most quickly. While DKM also converges very quickly, we point out in Table II that the final performance is much poorer.
@@ V Results
DKM 0.3613 0.8724 1.7685 0.1994
Ours + Formulation 1 0.3880 0.8241 1.6783 0.1834
Ours + Formulation 2 0.3892 0.8217 1.6710 0.1829
Ours + Formulation 3 0.3914 0.8102 1.6634 0.1792
Ours + Formulation 4 0.3910 0.8220 1.6782 0.1830
@@ V-A Performance on Waymo Motion Prediction Dataset
We show results for the Waymo Motion dataset in Table I, where we compared performance across two baselines, MTR [7] and DKM [1], and all formulations. From these results, we observe the greatest improvement over the baseline with Formulation 3, which involves the first-order velocity and heading components. We observe that the benefit of our method in full-scale training settings diminishes. This performance gap closing may be due to the computational complexity of the network, the large dataset, and or the long supervised training time out-scaling benefits provided by applying kinematic constraints. However, deploying models in the wild may not necessarily have such optimal settings, especially in cases of sparse data or domain transfer. Thus, we also consider the effects of suboptimal settings for trajectory prediction, as we hypothesize that learning first or second-order terms provides information when data cannot. This is also motivated by problems in the real world, where sensors may not be as high-quality or specific traffic scenarios may not be so abundantly represented in data.
TABLE II: Performance comparison on vehicles for each kinematic formulation versus SOTA in a small dataset setting. We train models on 1% of the original Waymo Motion Dataset and use the same full evaluation set as that in Table I. We see pronounced improvements in performance metrics in settings with significantly less data available, with a nearly 13% mAP performance gain over the baseline and nearly 50% mAP performance gain compared to deterministic kinematic method from DKM [1] across most formulations.
@@
DKM 0.1283 1.6632 3.7148 0.4361
Ours + Formulation 1 0.1920 1.3053 2.7907 0.3501
Ours + Formulation 2 0.1860 1.4386 2.7788 0.3500
Ours + Formulation 3 0.1856 1.3322 2.7717 0.3470
Ours + Formulation 4 0.1846 1.2948 2.8291 0.3541
@@ V-B Performance on a Smaller Dataset Setting
We examine the effects of kinematic priors on a smaller dataset size. This is motivated by the fact driving datasets naturally have an imbalance of scenarios, where many samples are representative of longitudinal straight-line driving or stationary movement, and much less are representative of extreme lateral movements such as U-turns. Thus, large and robust benchmarks like the Waymo, Nuscenes, and Argoverse datasets are necessary for learning good models. However, large datasets are not always accessible depending on the setting. For example, traffic laws, road design, and natural dynamics vary by region. It would be infeasible to expect the same scale and robustness of data from every scenario in the world, and thus trajectory forecasting will run into settings with less data available.
We train the baseline model and all formulations on only 1% of the original Waymo dataset and benchmark their performance on 100% of the evaluation set in Table II. All experiments were trained over 50 epochs. In the small dataset setting, we observe that providing a kinematic prior in any form improves performance for minimum final displacement error (minFDE). Additionally, we observe better performance across all metrics for formulations 1, 2, and 1 with interpolation. Figure 4 shows how all kinematic formulations improve convergence speed over the baseline, with Formulation 4 (acceleration and steering) converging most quickly. Overall, Formulation 1 provides the greatest boost in mAP performance, with over $13\%$ gain over the vanilla baseline and $50\%$ over the deterministic baseline DKM, while Formulations 3 and 4 provide the greatest boost in distance metrics. In general, all formulations provide similar benefits in the small dataset setting, with the most general performance benefit coming from Formulation 3. While Formulation 4 also provides a comparable performance boost, the difference from Formulation 3 may be attributed to compounding error from second-order approximation. Interestingly, we note that modeling mean trajectories with kinematic priors is not enough to produce performance gain, as shown by the lower performance of DKM compared to vanilla MTR.
Compared to the results from Table I, the effects of kinematic priors in learning are more pronounced. Since kinematic priors analytically relate the position at one timestep to the position at the next, improvements in metrics may suggest that baseline models utilize a large amount of expressivity to model underlying kinematics. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can help to compensate by modeling simple constraints.
@@ V-C Performance in the Presence of Noise
We also show how kinematic priors can influence performance in the presence of noise. This is inspired by the scenario where sensors may have a small degree of noise associated with measurements dependent on various factors, such as weather, quality, interference, etc.
We evaluate the models from Table III when input trajectories are perturbed by standard normal noise $n_{\epsilon}\sim\mathcal{N}(0,1)$; results for performance degradation are shown in Table III.
We compute results in Table III by measuring the % of degradation of the perturbed evaluation from the corresponding original clean evaluation. We find that Formulation 4 from Section IV-B4 with steering and acceleration components preserves the most performance in the presence of noise. This may be due to that second-order terms like acceleration are less influenced by perturbations on position, in addition to providing explicit bicycle-like constraints on vehicle movement. Additionally, distributions of acceleration are typically centered around zero regardless of how positions are distributed [33], which may provide more stability for learning.
@@
Ours + Formulation 1 -5.271 1.436 0.710 1.849
Ours + Formulation 2 -5.780 1.809 0.897 2.475
Ours + Formulation 3 -3.445 1.052 0.369 1.510
Ours + Formulation 4 -2.028 1.165 1.375 0.823
@@ VI Discussion and Conclusion
In this paper, we present a simple method for including kinematic relationships in probabilistic trajectory forecasting. Kinematic priors can also be implemented for deterministic methods where linear approximations are not necessary. With nearly no additional overhead, we not only show improvement in models trained on robust datasets but also in suboptimal settings with small datasets and noisy trajectories, with up to $12\%$ improvement in smaller datasets and $1\%$ less performance degradation in the presence of noise for the full Waymo dataset. For overall performance improvement, we find Formulation 1 with velocity components to be the most beneficial and well-rounded to prediction performance.

### v4

@@ IV-B1 Formulation 1: Velocity Components
Figure 4: Average Displacement Error (ADE) across training across each method and formulation in the small dataset setting. Learning with stochastic kinematic priors aids with faster learning. Compared to the baseline (blue), all models employing kinematic priors converge much more quickly. Formulation 4 (brown) converges most quickly.
@@ V Results
DKMUniform 0.3613 0.8724 1.7685 0.1994
DKMLearnable 0.3821 0.8150 1.6788 0.1839
OursF1 0.3880 0.8241 1.6783 0.1834
OursF2 0.3892 0.8217 1.6710 0.1829
OursF3 0.3914 0.8102 1.6634 0.1792
OursF4 0.3910 0.8220 1.6782 0.1830
@@ V-A Performance on Waymo Motion Prediction Dataset
We show results for the Waymo Motion dataset in Table I, where we compared performance across two baselines, MTR [7] and two variants of DKM [1], one with uniform variance and one with learnable variance (used in Multipath [29]), and all formulations. From these results, we observe the greatest improvement over the baseline with Formulation 3, which involves the first-order velocity and heading components. We observe that the benefit of our method in full-scale training settings diminishes. This performance gap closing may be due to the computational complexity of the network, the large dataset, and or the long supervised training time out-scaling benefits provided by applying kinematic constraints. However, deploying models in the wild may not necessarily have such optimal settings, especially in cases of sparse data or domain transfer. Thus, we also consider the effects of suboptimal settings for trajectory prediction, as we hypothesize that learning first or second-order terms provides information when data cannot.
TABLE II: Performance comparison on vehicles for each kinematic formulation versus SOTA in a small dataset setting. We train models on 1% of the Waymo Motion Dataset and use the same full evaluation set as that in Table I. We see pronounced improvements in performance metrics in settings with significantly less data available, with a nearly 13% mAP performance gain over the baseline and nearly 50% mAP performance gain compared to deterministic kinematic method from DKM [1] across most formulations. In the case of Formulation 4, analytical modeling of variance contributes an additional 2.5% in mAP performance over DKM with Learnable Covariance (DKMLearnable), which is employed by MultiPath [29].
@@
DKMUniform 0.1283 1.6632 3.7148 0.4361
DKMLearnable 0.1811 1.3082 2.9044 0.3612
OursF1 0.1907 1.2819 2.7956 0.3489
OursF2 0.1932 1.4141 2.6953 0.3376
OursF3 0.1833 1.3525 2.8362 0.3470
OursF4 0.1859 1.2932 2.8271 0.3502
@@ V-B Performance on a Smaller Dataset Setting
We examine the effects of kinematic priors on a smaller dataset size. This is motivated by imbalance of scenarios in driving datasets, where many samples are representative of longitudinal straight-line driving or stationary movement, and much less are representative of extreme lateral movements such as U-turns. Thus, large and robust benchmarks like the Waymo, Nuscenes, and Argoverse datasets are necessary for learning robust models.
We train the baseline model and all formulations on only 1% of the original Waymo dataset and benchmark their performance on 100% of the evaluation set in Table II. All experiments were trained over 50 epochs. In the small dataset setting, we observe that providing a kinematic prior in any form improves performance on all metrics except the deterministic case of DKM. Figure 4 shows how all kinematic formulations improve convergence speed over the baseline, with Formulation 4 (acceleration and steering) converging most quickly. Overall, Formulation 2 provides the greatest boost in mAP performance in the small dataset setting, with over $13\%$ mAP gain over the vanilla baseline and $50\%$ mAP gain over the deterministic baseline DKM. In general, all formulations provide similar benefits in the small dataset setting, with the most general performance benefit coming from Formulation 2. While Formulation 4 also provides a comparable performance boost, the performance difference from Formulation 1 and 2 may be attributed to compounding error from second-order approximation. Interestingly, we note that modeling mean trajectories with kinematic priors is not enough to produce performance gain, as shown by the lower performance of DKM, which is deterministic with unit variance, and Formulation 4 with Learnable Covariance (DKMLearnable in Table II).
Compared to the results from Table I, the effects of kinematic priors in learning are more pronounced in smaller dataset settings. Since kinematic priors analytically relate the position at one timestep to the position at the next, improvements in metrics may suggest that baseline models utilize a large amount of expressivity to model underlying kinematics. In backpropagation, optimization of one position further into the time horizon directly influences predicted positions at earlier timesteps via the kinematic model. Without the kinematic prior, the relation between timesteps may be implicitly related through neural network parameters. When the model lacks data to form a good model of how an agent moves through space, the kinematic model can help to compensate by modeling simple constraints.
@@ V-C Performance in the Presence of Noise
We also show how kinematic priors can influence performance in the presence of noise. This is inspired by the scenario where sensors may have a small degree of noise associated with measurements dependent on various factors, such as weather, quality, interference, etc. or compounding error from other modules in the autonomy stack.
We evaluate the models from Table I when input trajectories are perturbed by standard normal noise $n_{\epsilon}\sim\mathcal{N}(0,1)$; results for performance degradation are shown in Table III.
Table III results are computed by measuring the % of degradation of the perturbed evaluation from the corresponding original clean evaluation. We find that Formulation 4 from Section IV-B4 with steering and acceleration components preserves the most performance in the presence of noise. This may be due to that second-order terms like acceleration are less influenced by perturbations on position, in addition to providing explicit constraints on vehicle movement. Additionally, distributions of acceleration are typically centered around zero regardless of how positions are distributed [33], which may provide more stability in learning.
@@
DKMLearnable -1.268 1.404 1.068 1.223
OursF1 -5.271 1.436 0.710 1.849
OursF2 -5.780 1.809 0.897 2.475
OursF3 -3.445 1.052 0.369 1.510
OursF4 -2.028 1.165 1.375 0.823
@@ VI Discussion and Conclusion
In this paper, we present a simple method for including kinematic relationships in probabilistic trajectory forecasting. Kinematic priors can also be implemented for deterministic methods where linear approximations are not necessary. With nearly no additional overhead, we not only show improvement in models trained on robust datasets but also in suboptimal settings with small datasets and noisy trajectories, with up to $12\%$ improvement in smaller datasets and $1\%$ less performance degradation in the presence of noise for the full Waymo dataset. We find Formulation 2 with acceleration components to achieve the best rote performance, but Formulation 4 to be the most robust and fastest to converge. All experiments with kinematics modeled achieve non-trivial improved performance over the baseline model predicting trajectories directly.
@@
Acknowledgement: This research is supported in part by IARPA HAYSTAC Program, Barry Mersky and Capital One E-Nnovate Endowed Professorships.